aneurysm forms in a small blood vessel through which blood travels at 3 m·s–1. The
diameter of the blood vessel increases by 20%. What is the increase in pressure inside this aneurysm?
(ρblood= 1060 kg·m–3.)
Answer #1

using the continuity equation we have

A1v1 = A2v2

pi/4*d1^2 * v1 = pi/4*d2^2 * v2

(d)^2 * 3 = (1.2d)^2 * v2

v2 = 2.083 m/s

Using Bernoulli's equation we have

P1 + 0.5pv1^2 = P2 + 0.5pv2^2

P2 - P1 = 0.5*p ( v1^2 - v2^2)

= 0.5*1060 * ( 3^2 - 2.083^2)

= 247 Pa

so increase in pressure is 247 Pa